Problems In th e problems below, you can use a market risk premium of 55% and a tax rate of 40% where none is specified. 2. You are analyzing the balance sheet for Bed, Bath and Beyond, a retail firm that sells home furnishings, from Febru- ary 26, 1995 (in millions): Liabilities Assets Accounts Payable Other Current Liabilities $27.5 18.6 Cash Receivables Inventory $6.5 3.1 108.4 118.0 53.8 Current Liabilities Long-Term Debt Equity Total Liabilities 46.1 16.8 108.9 $171.8 Current Assets Fixed Assets Total Assets $171.8 The firm had revenues of $440.3 million in 1994 and cost of goods sold of $249.2 million. a. Estimate the net working capital. b. Estimate the noncash working capital c. Estimate the noncash working capital as a percent cash working capital needs for a new store for Bed, Bath and Beyond, would you use this ratio? Why or why not? of revenues. If you were asked to estimate the non-
Answer #1

a) Net Working Capital

Net Working Capital= Current Assets - Current Liabilities

Current Liability= $ 46.1 Million

Current Assets =$ 118 Million

Net Working Capital= 118-46.1

= $ 79.1 Million

b) Non Cash Working = Current Assets excluding Cash - Current Liabilities

Current Assets excluding Cash= $ 118- $ 6.5

=$111.5 Million

Non Cash Working Capital= 111.5- 46.1

= $ 65.4 Million

c) Non cash working capital as a percentage of revenue

Non Cash Working Capital= $ 65.4 Million

Revenue in 1994= $ 440.3 Million

Percentage = 65.4/440.3*100

14.85%

For new Store i would use this ration since it will provide the better Working capital to turnover ratio and give high number of working capital cycle hence working capital utilization is high.